OpenStudy (amtran_bus):

Derivative of 3x^2 - 2 cos x

OpenStudy (amtran_bus):

Well, Is it 6x+2 sin x or do I need chain rule?

OpenStudy (jacob902):

Distribute the 3. 3x + 3 + 6 = 33 Isolate the x term. 3x = 33 - 3 - 6 Simplify. 3x = 24 Divide. x = 8 Check answer. 3[(8) + 1] + 6 = 33 3[9] + 6 = 33 27 + 6 = 33 33 = 33

OpenStudy (pratyush5):

You have the correct answer

OpenStudy (amtran_bus):

Why do I not do it this way? Let me type